OpenStudy (anonymous):

-5x-y=7 5x+y=2 solve system of equation. can someone help me please!?

sam (.sam.):

You can solve by either substitution or elimination

sam (.sam.):

-5x-y=7.....(1) 5x+y=2.....(2) -------------- Take (2), 5x+y=2 y=2-5x.....(3) -------------- Substitute (3) into (1), -5x-(2-5x)=7 You can find x from here

OpenStudy (anonymous):

not likely

OpenStudy (anonymous):

i'm lost!

OpenStudy (anonymous):

multiply the first equation by \(-1\) and get an equivalent equation \[5x+y=-7\]

OpenStudy (anonymous):

so you have two equations \(a)\) \(5x+y=-7\) \(b)\) \(5x+y=2\)

OpenStudy (anonymous):

it is not possible for \(5x+y\) to be both \(-7\) and \(2\) so you cannot solve this the equations are for two lines they are parallel, so they will not intersect

OpenStudy (anonymous):

These two equations represent parallel lines for which there is no common point. There is no solution to this set of equations. That should be apparent from the start. Add the left hand sides together and you get zero on the left. Add the right hand sides together and you get 9 over there. Sorry, but zero does not equal nine!

OpenStudy (anonymous):

thanks